OpenStudy (anonymous):

Which of the following is not a step used when graphing the equation y = −x + 8 using the slope-intercept method? Plot the x-intercept (32, 0). Plot the y-intercept (0, 8). Draw a line through the two points. Count down 1 and right 4 from the y-intercept.

OpenStudy (anonymous):

What do you think tell me first..

OpenStudy (anonymous):

Any reason?? Or just guessing??

OpenStudy (anonymous):

im not guessin

OpenStudy (anonymous):

See can you tell me what is x intercept here??

OpenStudy (anonymous):

Just put y = 0 as you did earlier...

OpenStudy (anonymous):

probably would not graph the x-intercept. The most logically thing to do would be to graph the y-intercept and then count down 1 (I'm going down 1 because of the negative sign) and go right 4. Another way to do it would be to create a "T" table, where you plug in values for x and evaluate y using -1/4x + 8. I prefer this method and I will use a minimum of three x values to plug in to make sure I am not making a mistake. Some other things to consider - whenever you have a NEGATIVE slope, the graph should be going down to the right, and a POSITIVE slope should be going up to the right.

OpenStudy (anonymous):

I think you know more than me but you have not faith on yourself.. Yes you are right again..

OpenStudy (anonymous):

lol so wat should i choose

OpenStudy (anonymous):

b righ

OpenStudy (anonymous):

A is the option for x intercept ?? Is it??

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

So what should be the option??

OpenStudy (anonymous):

I think first option ( 32,0) Is not the step in the plotting of this line on graph

OpenStudy (anonymous):

It is asking to graph using the slope-intercept method. THis method involves obtaining the y-intercept and then following the slope. This means plot the y-intercept and then plot another point using the slope and then draw the line through it. Therefore, fo r this particular method, there is no need for an x-intercept. The fact that the form y = mx + b has y-intercept and slope, but not the x-intercept further proves the point. That being said, it is choice ?

OpenStudy (anonymous):

@Diana97 make sure that from next time you type the question correctly.. y = -x + 8 Incorrect.. y = (-1/4)x + 8 is the original equation...

OpenStudy (anonymous):

i guess is d

OpenStudy (anonymous):

I said that there is no need for an x-intercept. Which answer choice involves the x-intercept?

OpenStudy (anonymous):

a